# Probing Cosmic Curvature with Fast Radio Bursts and DESI DR2

### v1

@@ VI Results and Discussions
To apply the methodology described above, we first subtract the contributions from the Milky Way, the halo, and the host galaxy in order to obtain $\mathrm{DM_{IGM}}$ from the observed dispersion measures of the subset of 80 FRBs described in Sec. V. Then we reconstruct the comoving distance $D_{C}(z)$. The first step involves determining the redshift evolution of the average intergalactic dispersion measure, $\langle\mathrm{DM}_{\mathrm{IGM}}(z)\rangle$. As detailed in Sec. IV, this is performed using an ANN algorithm (see Fortunato et al. (2025) for more details). The result is shown in panel (a) in Figure 1. The red points with error bars represent the individual FRB measurements and their associated uncertainties, while the solid black curve corresponds to the ANN-based reconstruction. The shaded region around it denotes the $1\sigma$ confidence interval.
@@
Figure 1: Reconstructed FRBs observables. Left: redshift evolution of the average IGM dispersion measure $\langle\mathrm{DM}_{\mathrm{IGM}}(z)\rangle$. Right: comoving distance $D_{C}(z)$ inferred from FRBs.
@@
Once $\langle\mathrm{DM}_{\mathrm{IGM}}(z)\rangle$ is reconstructed, it is used to derive the comoving distance $D_{C}(z)$ (see Eq. (17)). The resulting comoving distance curve is displayed in panel (b) in Figure 1. The dashed black line shows the reconstructed $D_{C}(z)$, while the shaded blue band represents the associated $1\sigma$ uncertainty. This reconstruction provides a continuous cosmological model-independent estimate of the comoving distance across a range of redshifts, from low redshifts up to $z\sim 1.5$, where BAO measurements are also available for comparison. Subsequently, $D^{\mathrm{FRB}}_{A}(z)$ and $D^{\mathrm{BAO+FRB}}_{A}(z)$ are computed.
@@
To constrain $\Omega_{k}$, we explore the likelihood using the emcee Python module, a Markov Chain Monte Carlo (MCMC) ensemble sampler Foreman-Mackey et al. (2013). The prior for the Hubble constant is taken as the extrapolated value $H_{0}=H(z=0)$, obtained from the ANN reconstruction of the FRB data. The final results from both approaches (covariance matrix and $\chi^{2}$ minimization) are presented and compared in Table 2, and the corresponding posterior distributions are shown in Figure 2. These results demonstrate that the combined use of FRBs and BAO data provides a robust and independent constraint on the spatial geometry of the Universe.
Figure 2: Posterior probability distributions for the cosmic curvature parameter $\Omega_{k}$ under two assumptions. Left: using the full covariance matrix Right: Neglecting correlations.
Survey $\Omega_{k}$ (with covariance) $\Omega_{k}$ (without covariance)
eBOSS $-0.25^{+0.35}_{-0.46}$ $-0.18\pm 0.20$
DESI DR2 $0.10^{+0.33}_{-0.38}$ $-0.19\pm 0.08$
eBOSS + DESI DR2 $-0.09^{+0.24}_{-0.31}$ $-0.17\pm 0.06$
Table 2: Estimated values of the cosmic curvature parameter $\Omega_{k}$ using BAO and FRB comoving distance measurements.
From the covariance-based analysis (Table 2), we derive the following constraints:
  • The eBOSS data yield $\Omega_{k}=-0.25^{+0.35}_{-0.46}$, indicating a mild preference for a closed Universe, albeit with substantial uncertainties.
  • The DESI data alone provide a constraint centered around spatial flatness, with $\Omega_{k}=0.10^{+0.33}_{-0.38}$.
  • When combining the eBOSS and DESI DR2 datasets, the inferred curvature is $\Omega_{k}=-0.09^{+0.24}_{-0.31}$, suggesting a slightly closed Universe with reduced uncertainties compared to individual datasets.
In contrast, neglecting the off-diagonal elements of the covariance matrix (Table 2) yields significantly tighter posterior distributions:
  • For eBOSS, we find $\Omega_{k}=-0.18\pm 0.20$,
  • For DESI, $\Omega_{k}=-0.19\pm 0.08$, and
  • For the joint eBOSS + DESI DR2 analysis, $\Omega_{k}=-0.17\pm 0.06$.
@@
Despite variations in central values, all estimates remain consistent with a spatially flat Universe ($\Omega_{k}=0$) at the $1\sigma$ level. The integration of FRB-based comoving distance reconstruction serves as an independent and complementary cosmological probe, reinforcing the growing utility of FRBs in precision cosmology.
Taken together, these results demonstrate that ANN reconstructions of comoving distances from FRBs, when combined with state-of-the-art BAO measurements, provide competitive constraints on spatial curvature, comparable to those obtained from other late-time cosmological probes Liu et al. (2025); Gong et al. (2024); Mukherjee and Banerjee (2022); Wang et al. (2021). While Gaussian-based methods yield tighter intervals, they may underestimate uncertainties due to neglected correlations. The apparent preference for negative values in this case may result from omitting the covariance matrix, which can bias the final result. In contrast, the approach that incorporates the full covariance matrix, though yielding broader confidence regions, provides a more conservative and robust assessment of parameter uncertainties. These findings highlight the critical role of accounting for the full covariance structure in cosmological parameter inference, especially when combining different datasets or when intrinsic correlations are present.
@@ VII Conclusions
In this work, we presented a fully data-driven strategy to constrain the spatial curvature of the Universe by leveraging the synergy between FRBs and BAO measurements. Reconstructing the Hubble parameter $H(z)$ using an ANN trained on a set of localized FRBs, we derived the comoving distance $D_{C}(z)$ and subsequently estimated $D_{A}(z)$ to compare with the same quantity obtained from BAO+FRB data. This dual-path approach enabled robust constraints on the curvature parameter $\Omega_{k}$, independent of early-universe assumptions or fiducial cosmological models.
@@ Acknowledgements.
JASF thanks FAPES for their financial support. WSHR acknowledges partial financial support from FAPES. G.E.R. acknowledges financial support from the State Agency for Research of the Spanish Ministry of Science and Innovation under grant PID2022-136828NB-C41/AEI/10.13039/501100011033/, and by “ERDF A way of making Europe”, by the “European Union”. He also thanks support from PIP 0554 (CONICET).

### v2

@@ VI Results and Discussions
To apply the methodology described above, we first subtract the contributions from the Milky Way, the halo, and the host galaxy in order to obtain $\mathrm{DM_{IGM}}$ from the observed dispersion measures of the subset of 120 FRBs described in Sec. V. Then we reconstruct the comoving distance $D_{C}(z)$. The first step involves determining the redshift evolution of the average intergalactic dispersion measure, $\langle\mathrm{DM}_{\mathrm{IGM}}(z)\rangle$. As detailed in Sec. IV, this is performed using an ANN algorithm (see Fortunato et al. (2025) for more details). The result is shown in panel (a) in Figure 1. The red points with error bars represent the individual FRB measurements and their associated uncertainties, while the solid black curve corresponds to the ANN-based reconstruction. The shaded region around it denotes the $3\sigma$ confidence interval.
@@
Figure 1: Reconstructed FRBs observables. Left: redshift evolution of the average IGM dispersion measure $\langle\mathrm{DM}_{\mathrm{IGM}}(z)\rangle$. Right: Hubble parameter $H(z)$ inferred from FRBs.
@@
Once $\langle\mathrm{DM}_{\mathrm{IGM}}(z)\rangle$ is reconstructed, it is used to derive the Hubble parameter $H(z)$ and by extrapolating the reconstruction to $z=0$ we obtain $H_{0}=68.74\pm 6.11\ \mathrm{km\,s^{-1}\,Mpc^{-1}}$. Panel (b) of Figure 1 shows the reconstructed $H(z)$ curve with $3\sigma$ bands. This value is then applied throughout our methodology as described in Section IV to recover the comoving distance $D_{C}(z)$ (see Eq. (18)). The resulting comoving distance curve is displayed in Figure 2. The dashed black line shows the reconstructed $D_{C}(z)$, while the shaded blue band represents the associated $3\sigma$ uncertainty. This reconstruction provides a continuous cosmological model-independent estimate of the comoving distance across a range of redshifts, from low redshifts up to $z\sim 2.5$, where BAO measurements are also available for comparison. Subsequently, $D^{\mathrm{FRB}}_{A}(z)$ and $D^{\mathrm{BAO+FRB}}_{A}(z)$ are computed.
Figure 2: Comoving distance $D_{C}(z)$ reconstructed from FRBs.
@@
To constrain $\Omega_{k}$, we explore the likelihood using the emcee Python module, a Markov Chain Monte Carlo (MCMC) ensemble sampler Foreman-Mackey et al. (2013). The prior for the Hubble constant is taken as the extrapolated value $H_{0}=H(z=0)$, obtained from the ANN reconstruction of the FRB data. The final results from both approaches (covariance matrix and Gaussian $\chi^{2}$ minimization) are presented in Figure 3, panel (a). These results demonstrate that the combined use of FRBs and BAO data provides a robust and independent constraint on the spatial geometry of the Universe.
Figure 3: Left: One–dimensional posterior for the spatial curvature parameter $\Omega_{k}$ obtained with our two likelihoods. The red filled curve (COV) uses the full covariance between the FRB–derived $D_{C}$ and $D_{A}$ together with the BAO error propagation; the blue filled curve (Gaussian) adopts a diagonal treatment. The vertical dashed line marks $\Omega_{k}=0$. Right: Black points with $3\sigma$ bars are $D_{A}^{\rm BAO+FRB}(z)$ from Eq. 14, using DESI DR2 data with the FRB–based $H(z)$ and propagated BAO uncertainties. Colored curves show the FRB–only estimator $D_{A}^{\rm FRB}(z;\Omega_{k})$ evaluated at the mean $\Omega_{k}$ of each analysis; the shaded bands denote credible regions.
@@
From the covariance-based analysis by employing our 120 FRBs and the DESI DR2 datasets, we derive an estimated $\Omega_{k}=-0.31\pm 0.57$ for the cosmic curvature parameter. In contrast, neglecting the off-diagonal elements of the covariance matrix yields a bit tighter posterior distribution, $\Omega_{k}=-0.13\pm 0.46$. Despite variations in central values, all estimates remain consistent with a spatially flat Universe ($\Omega_{k}=0$) at the $1\sigma$ level. The integration of FRB-based comoving distance reconstruction serves as an independent and complementary cosmological probe, reinforcing the growing utility of FRBs in precision cosmology.
In panel (b) of Figure 3 we compare the FRB-only reconstruction of the angular–diameter distance, $D_{A}^{\rm FRB}(z;\Omega_{k})$, obtained with our two likelihood treatments against the discrete $D_{A}^{\rm BAO+FRB}$ estimates at the DESI DR2 effective redshifts – black squares. Solid curves show the posterior–mean $D_{A}^{\rm FRB}(z)$ for each treatment, while the shaded bands indicate the propagated $3\sigma$ uncertainty, which naturally widens toward high $z$ where the FRB sample is sparse. The DESI points (with $1\sigma$ error bars) are derived from the measured BAO ratios via Eq. 14 combined with the FRB-based $H(z)$ reconstruction. The two FRB-only solutions agree closely over $0.4\!\lesssim\!z\!\lesssim\!2$, with the COV band being broader—as expected when correlations are retained—yet both envelopes encompass the BAO+FRB determinations within uncertainties. The peak around $z\!\sim\!1.6$ is consistent with the expected maximum of $D_{A}(z)$ in late-time cosmologies, and the overall consistency of the three estimators supports our curvature analysis.
Taken together, these results demonstrate that ANN reconstructions of comoving distances from FRBs, when combined with state-of-the-art BAO measurements, provide competitive constraints on spatial curvature, comparable to those obtained from other late-time cosmological probes Liu et al. (2025); Gong et al. (2024); Mukherjee and Banerjee (2022); Wang et al. (2021). In our analysis, both likelihood treatments are statistically consistent with spatial flatness at the $1\sigma$ level, but they behave differently with respect to uncertainty budgeting. The Gaussian approximation produces visibly tighter credible intervals, yet it risks understating the true errors because it neglects cross–terms induced by the common FRB–based $H(z)$ entering both $D_{A}$ and $D_{C}$. When we retain the full covariance, the posteriors broaden as expected, yielding a more conservative and robust assessment of parameter uncertainties. Importantly, the mild preference for $\Omega_{k}<0$ persists in both treatments, and the shift of the posterior mean when toggling the covariance is small, indicating that our qualitative conclusion is not driven by the covariance choice.
Figure 4: Constraints on the spatial curvature $\Omega_{k}$. Horizontal bars show $1\sigma$ confidence intervals for our FRB + BAO (COV) and FRB + BAO (GAUSS) analyses, alongside a set of literature determinations (Planck 2018 + BAO; Planck 2018 (no lensing); EFTofLSS/BAO; BAO + Pantheon+; CC + SGL; Planck 2018 + $C$; etc.). The vertical dashed line marks $\Omega_{k}=0$. Incorporating the full covariance (COV) yields broader, more conservative constraints than the Gaussian approximation (GAUSS), yet both remain consistent with spatial flatness at $1\sigma$.
@@ VII Conclusions
In this work, we presented a fully data-driven strategy to constrain the spatial curvature of the Universe by leveraging the synergy between FRBs and BAO measurements. Reconstructing the Hubble parameter $H(z)$ using an ANN trained on a set of localized FRBs, we derived the comoving distance $D_{C}(z)$ and subsequently estimated $D_{A}(z)$ to compare with the same quantity obtained from BAO+FRB data. This dual-path approach enabled robust constraints on the curvature parameter $\Omega_{k}$, independent of early-universe assumptions or fiducial cosmological models. Our analysis yields $\Omega_{k}=-0.31\pm 0.57$ (full covariance) and $\Omega_{k}=-0.13\pm 0.46$ (Gaussian approximation). While the uncorrelated treatment produces tighter intervals, it risks underestimating uncertainties, we therefore adopt the covariance-based constraint as our headline result.
@@ Acknowledgements.
JASF thanks FAPES and the National Research Foundation (NRF) for their financial support. WSHR acknowledges partial financial support from FAPES. G.E.R. acknowledges financial support from the State Agency for Research of the Spanish Ministry of Science and Innovation under grant PID2022-136828NB-C41/AEI/10.13039/501100011033/, and by “ERDF A way of making Europe”, by the “European Union”. He also thanks the support from PIP 0554 (CONICET). JASF thanks Amanda Weltman and Surajit Kalita for fruitful discussions.
.1 Dataset
In this appendix we present the Full Localized FRBs sample, including the ones used in this work and those we neglected for the reasons described in the main text. Reported observables are the measured dispersion measure $\mathrm{DM}_{\rm obs}$ and the Milky Way contribution from the NE2001 model.
Table 2: Localized FRBs
FRB $z$ $\mathrm{DM}_{\rm obs}$ $\mathrm{DM}_{\rm MW}$ (NE2001) Ref.
FRB200120E 0.0008 87.8 40.8 Connor et al. (2024)
FRB181030A 0.00385 103.396 41.1 Lanman et al. (2022)
FRB171020A 0.008672 114.1 38.0 Mahony et al. (2018)
FRB220319D 0.011228 110.98 65.25 Ravi et al. (2023)
FRB231129A 0.019000 198.50 58.3 Amiri et al. (2025)
FRB240210A 0.023686 283.73 28.7 Shannon et al. (2025)
FRB181220A 0.027460 208.66 118.5 Bhardwaj et al. (2024)
FRB231230A 0.029800 131.40 23.0 Amiri et al. (2025)
FRB181223C 0.030240 111.61 19.9 Bhardwaj et al. (2024)
FRB190425A 0.03122 127.78 48.7 Bhardwaj et al. (2024)
FRB180916B 0.033700 349.35 199.0 Marcote et al. (2020)
FRB230718A 0.0357 477.0 420.6 Glowacki et al. (2024)
FRB231120A 0.036800 437.74 43.8 Sharma et al. (2024)
FRB240201A 0.042729 374.50 38.6 Shannon et al. (2025)
FRB220207C 0.043040 262.38 76.1 Law et al. (2024)
FRB211127I 0.046900 234.83 42.5 Glowacki et al. (2023)
FRB201123A 0.050700 433.55 251.7 Rajwade et al. (2022)
FRB230926A 0.055300 222.80 52.6 Amiri et al. (2025)
FRB200223B 0.060240 201.80 45.6 Ibik et al. (2024)
FRB190303A 0.064000 223.20 29.8 Michilli et al. (2023)
FRB231204A 0.064400 221.00 34.9 Amiri et al. (2025)
FRB231206A 0.065900 457.70 59.1 Amiri et al. (2025)
FRB210405I 0.066 565.17 516.1 Driessen et al. (2024)
FRB180814A 0.068 190.9 87.5 Michilli et al. (2023)
FRB211212A 0.070700 206.00 38.8 et al.
FRB231005A 0.071300 189.40 33.5 Amiri et al. (2025)
FRB190418A 0.071320 182.78 70.2 Bhardwaj et al. (2024)
FRB231123A 0.072900 302.10 89.7 Amiri et al. (2025)
FRB220912A 0.0771 219.46 125.2 Ravi et al. (2023)
FRB231011A 0.078300 186.30 70.4 Amiri et al. (2025)
FRB220509G 0.089400 269.53 55.6 Law et al. (2024)
FRB230124A 0.093900 590.57 38.6 Sharma et al. (2024)
FRB201124A 0.098000 413.52 139.9 Ravi et al. (2022)
FRB230708A 0.105000 411.51 60.3 Shannon et al. (2025)
FRB231223C 0.105900 165.80 47.9 Amiri et al. (2025)
FRB191106C 0.107750 332.20 25.0 Ibik et al. (2024)
FRB231128A 0.107900 331.60 64.7 Amiri et al. (2025)
FRB230222B 0.110000 187.80 27.7 Amiri et al. (2025)
FRB231201A 0.111900 169.40 70.0 Amiri et al. (2025)
FRB220914A 0.113900 631.28 54.7 Law et al. (2024)
FRB190608B 0.117780 338.70 37.3 Chittidi et al. (2021)
FRB230703A 0.118400 291.30 26.9 Amiri et al. (2025)
FRB240213A 0.118500 357.40 40.0 Connor et al. (2024)
FRB230222A 0.122300 706.10 134.2 Amiri et al. (2025)
FRB190110C 0.122440 221.60 37.1 Ibik et al. (2024)
FRB230628A 0.127000 344.95 39.0 Sharma et al. (2024)
FRB240310A 0.127000 601.80 30.1 Shannon et al. (2025)
FRB210807D 0.129300 251.90 121.2 Glowacki et al. (2023)
FRB240114A 0.130000 527.65 49.7 Kumar et al. (2024)
FRB240209A 0.138400 176.57 55.5 Shah et al. (2025)
FRB210410D 0.141500 571.20 56.2 Caleb et al. (2023)
FRB230203A 0.146400 420.10 67.3 Amiri et al. (2025)
FRB231226A 0.156900 329.90 38.1 Shannon et al. (2025)
FRB230526A 0.157000 361.40 31.9 Shannon et al. (2025)
FRB220920A 0.158239 314.99 39.9 Law et al. (2024)
FRB200430A 0.160800 380.10 27.2 Heintz et al. (2020)
FRB190203A 0.170000 134.40 20.3 Tyul’bashev et al. (2025)
FRB210603A 0.177200 500.15 39.5 Cassanelli et al. (2024)
FRB220529A 0.183900 246.00 40.0 Li et al. (2025)
FRB230311A 0.191800 364.30 92.5 Amiri et al. (2025)
FRB220725A 0.192600 290.40 30.7 Shannon et al. (2025)
FRB121102A 0.192730 557.00 188.4 Tendulkar et al. (2017)
FRB221106A 0.204400 343.80 34.8 Shannon et al. (2025)
FRB240215A 0.210000 549.50 47.9 Connor et al. (2024)
FRB230730A 0.211500 312.50 85.2 Amiri et al. (2025)
FRB210117A 0.214500 729.10 34.4 Bhandari et al. (2023)
FRB221027A 0.229000 452.50 47.2 Sharma et al. (2024)
FRB191001A 0.234000 506.92 44.2 Heintz et al. (2020)
FRB190714A 0.236500 504.13 38.5 Heintz et al. (2020)
FRB221101B 0.239500 491.55 131.2 Sharma et al. (2024)
FRB220825A 0.241397 651.24 78.5 Law et al. (2024)
FRB190520B 0.2418 1204.7 63 Ocker et al. (2022)
FRB191228A 0.243200 297.50 32.9 Bhandari et al. (2022)
FRB231017A 0.245000 344.20 31.2 Amiri et al. (2025)
FRB220307B 0.248123 499.27 128.2 Law et al. (2024)
FRB221113A 0.250500 411.03 91.7 Sharma et al. (2024)
FRB220831A 0.262 1146.25 126.8 Connor et al. (2024)
FRB231123B 0.262100 396.86 40.3 Amiri et al. (2025)
FRB230307A 0.270600 608.85 37.6 Sharma et al. (2024)
FRB221116A 0.276400 643.45 132.3 Sharma et al. (2024)
FRB220105A 0.278500 583.00 22.0 Shannon et al. (2025)
FRB210320C 0.279700 384.80 39.3 Glowacki et al. (2023)
FRB221012A 0.284669 441.08 54.3 Law et al. (2024)
FRB240229A 0.287000 491.15 38.0 Connor et al. (2024)
FRB190102C 0.291200 364.50 57.4 Macquart et al. (2020)
FRB220506D 0.300390 396.97 84.6 Law et al. (2024)
FRB230501A 0.301500 532.47 125.7 Sharma et al. (2024)
FRB230503E 0.320000 483.74 88.0 Pastor-Marazuela et al. (2025)
FRB180924B 0.321200 361.42 40.5 Bannister et al. (2019)
FRB231025B 0.323800 368.70 48.6 Amiri et al. (2025)
FRB230125D 0.3265 640.08 88.0 Pastor-Marazuela et al. (2025)
FRB230626A 0.327 452.723 39.3 Sharma et al. (2024)
FRB180301A 0.330400 552.00 151.7 Price et al. (2019)
FRB231220A 0.335500 491.20 49.9 Connor et al. (2024)
FRB211203C 0.343900 636.20 63.7 Shannon et al. (2025)
FRB230808F 0.347200 653.20 27.0 Hanmer et al. (2025)
FRB220208A 0.351000 437.00 101.6 Sharma et al. (2024)
FRB230902A 0.361900 440.10 34.1 Shannon et al. (2025)
FRB220726A 0.361900 686.23 89.5 Sharma et al. (2024)
FRB220717A 0.362950 637.34 118.3 Rajwade et al. (2024)
FRB200906A 0.3688 577.8 35.8 Bhandari et al. (2022)
FRB220330D 0.3714 468.1 38.6 Sharma et al. (2024)
FRB240119A 0.376000 483.10 38.0 Connor et al. (2024)
FRB190611B 0.377800 321.40 57.8 Heintz et al. (2020)
FRB220501C 0.381000 449.50 30.6 Shannon et al. (2025)
FRB230613A 0.392300 483.51 30.0 Pastor-Marazuela et al. (2025)
FRB220204A 0.401200 612.58 50.7 Connor et al. (2025)
FRB230712A 0.452500 587.57 39.2 Sharma et al. (2024)
FRB230907D 0.463800 1030.79 29.0 Pastor-Marazuela et al. (2025)
FRB181112A 0.475500 589.27 102.0 Prochaska et al. (2019)
FRB231020B 0.477500 952.20 34.0 Pastor-Marazuela et al. (2025)
FRB220307B 0.248123 499.27 128.2 Law et al. (2024)
FRB220310F 0.477958 462.24 46.3 Law et al. (2024)
FRB220918A 0.491000 656.80 41.0 Shannon et al. (2025)
FRB231210F 0.500000 720.60 32.0 Pastor-Marazuela et al. (2025)
FRB190711A 0.522000 593.10 56.5 Heintz et al. (2020)
FRB230216A 0.531000 828.00 38.5 Sharma et al. (2024)
FRB221219A 0.553000 706.71 44.4 Sharma et al. (2024)
FRB230814A 0.553000 696.40 104.8 Connor et al. (2024)
FRB190614D 0.600000 959.20 87.8 Law et al. (2024)
FRB210924D 0.600000 737.00 45.0 Carli et al. (2024)
FRB231010A 0.610000 442.59 41.0 Pastor-Marazuela et al. (2025)
FRB220418A 0.622000 623.25 36.7 Law et al. (2024)
FRB220224C 0.627100 1140.20 52.0 Sharma et al. (2024)
FRB190523A 0.660000 760.80 37.2 Ravi et al. (2019)
FRB190208A 0.830000 580.01 71.5 Hewitt et al. (2024)
FRB220222C 0.853000 1071.20 56.0 Pastor-Marazuela et al. (2025)
FRB240123A 0.968000 1462.00 90.2 Connor et al. (2024)
FRB221029A 0.975000 1391.75 43.8 Sharma et al. (2024)
FRB220610A 1.016000 1458.15 31.0 Gordon et al. (2024)
FRB230521B 1.354000 1342.90 138.8 Connor et al. (2024)
FRB240304B 2.148000 2330.00 28.1 Caleb et al. (2025)